OpenStudy (anonymous):

The length of the Boston Marathon is 138,435 ft. Each stride, or step, by a particular runner is 3ft long. Does the number of strides the runner takes fit evenly into the length of the race? Explain

OpenStudy (alekos):

Just add up all the individual digits and if the result is divisible by 3 then the original number is also.

OpenStudy (karolynxx):

By evenly the problem means that if you layed all the strides side by side, then the last one would end exactly on the finish line. Or more simply, if you divide the course length by the stride, the Remainder would be zero. There are no even or odd groups to worry about. So 138,435 / 3 = 46,145 strides exactly, so the strides fit evenly into the course. I could have a stride of 3.2 ft and win the race even though 138,435 / 3.2 = 43,260.9375 doesn't fit evenly. Then my last stride would land beyond the finish line and not just on it.

OpenStudy (anonymous):

can you help with more? @karolynxx

OpenStudy (osgood):

What else maybe I can help @rainbowswirl

OpenStudy (anonymous):

I'm still confused @karolynxx

OpenStudy (anonymous):

ok thanks @Osgood

OpenStudy (anonymous):

A number is written with the following factorization: 2 to the 2nd power x 3 x 5 to the 4th power x 8 x 11 to the 2nd power. Is this factorization a prime factorization. Explain why or why not. If it is not correct, give the correct prime factorization of the number. @Osgood @karolynxx

OpenStudy (osgood):

@xMissAlyCatx

OpenStudy (osgood):

IDK @rainbowswirl but @xMissAlyCatx will probably know I have to go

OpenStudy (anonymous):

oh, okay. I really need help @Osgood @xMissAlyCatx

OpenStudy (xmissalycatx):

What @karolynxx said was pretty solid, though.

OpenStudy (anonymous):

Did you see my other question? I figured out the one @karolynxx did. Could you help me? @xMissAlyCatx

OpenStudy (karolynxx):

I could have a stride of 3.2 ft and win the race even though 138,435 / 3.2 = 43,260.9375 doesn't fit evenly. Then my last stride would land beyond the finish line and not just on it i put that on there

OpenStudy (anonymous):

I dont get it @xMissAlyCatx

OpenStudy (karolynxx):

i see what your asking

OpenStudy (karolynxx):

The sum of the digits is: 24. This, the number is divisible by 3. Additionally, we know 120,000 is divisible by 3, 18,000 is divisible by 3, 420 is divisible by 3, and 15 is divisible by 3. so 3

OpenStudy (anonymous):

I think I got that one down @karolynxx. could you help with more though?

OpenStudy (anonymous):

Is it a prime factorization though? @karolynxx

OpenStudy (osgood):

Hey @rainbowswirl This site will describe prime factorization http://www.mathsisfun.com/prime-factorization.html